P(x)=2x^3-21x^2-2x+25; a=1,b=2
Found 2 solutions by solver91311, hasma812:
Answer by solver91311(24713)

What is it you want to do with this function? How do a and b relate? Are they supposed to be x-coordinates of the x-intercepts? Are you just supposed to evaluate the function at a and b? Give us something to go on here, please. This note applies to the other two problems that you submitted as well.

Answer by hasma812(1)

P(x)=2x^3-21x^2-2x+25; a=1,b=2
P(1)=2(1)^3-21(1)^2-2(1)+25
2-21-2+25
2-23+25
=4
P(2)=2(2)^3-21(2)^2-2(2)+25
2(8)-21(4)-4+25
16-84-4+25
-68-4+25
-72+25
=-47